OpenStudy (anonymous):

if i am not given the value of the first value in a problem, how do i find the y intercept?

OpenStudy (anonymous):

What do you mean?

OpenStudy (anonymous):

well, i am given 1.0 bushels, in billions, at a price of 1.35. each value of bushels increases by .5 and each relative price increases by a 1.05. this goes for 4 values up to 2.5 (billion) bushels at an ending price of 4.50. i have the slope of 2.1 but i need the y intercept or where x=0

OpenStudy (anonymous):

Wait, was the slope given or did the problem describe how much the price went up too?

OpenStudy (anonymous):

no, i was given a table with values. then i had to calculate the average rate of change(slope)

OpenStudy (anonymous):

for each price, the value is constant

OpenStudy (anonymous):

im scared you're typing for so long lol

OpenStudy (anonymous):

Ok, yes. So you're right. Your slope is 2.1 for this problem. So the easiest thing to do is start with the form y=mx+b. We know that m=2.1 and we're given that 1.0 bushels equals 1.35 in price. So we can plug in 1.0 bushels for y, 1.35 for x, 2.1 for m and then solve for b. 1.0=2.1(1.35)+b 1.0=2.835+b b=1.0-2.835 b=-1.835 I'm assuming price is plotted on the x-axis since price should start above 0 initially. So at this point, b is our y-intercept, so the coordinate pair (0, -1.835) is where your y-intercept is.

OpenStudy (anonymous):

Lol sorry I typed so long. I usually try to think how to explain things easily before I post. Its not so easy to explain the jumbled mess in your head.

OpenStudy (anonymous):

right, i appreciate the help!

OpenStudy (anonymous):

Make sense?

OpenStudy (anonymous):

but for this kind of problem you wouldn't necessarily want an x value of zero because money typically isnt negative.

OpenStudy (anonymous):

Right! If the price started at zero it would initially be free and that doesn't make sense! So, this usually means our y-intercept is negative and slope is positive when we talk about prices per bushel, in your problem, for instance.

OpenStudy (anonymous):

ok, so how can i test my formula?

OpenStudy (anonymous):

The y-axis is your independent value and the x-axis is the dependent value. Here, you would want bushels to be on the y-axis since price depends on the amount of bushels. So in the form y=mx+b, your equation is y=2.1x-1.835. To test, look at your table of values. Plug in a y-value (number of bushels) into the equation and solve. If your answer matches the corresponding price in the table, then your formula works!

OpenStudy (anonymous):

ummm i plugged in 1.5 bushels. 2.1*1.5-1.835=1.315 but the table at 1.5 says 2.40???

OpenStudy (anonymous):

sorry if im wasting your time

OpenStudy (anonymous):

Can you type the values in please? Maybe if I see them we can modify the equation to work. It appears we haven't done this right.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

1 bushel=1.35 1.5bushels=2.40 2bushels=3.45 4bushels=4.50

OpenStudy (anonymous):

Hold on, I got something backwards

OpenStudy (anonymous):

alright no problem

OpenStudy (anonymous):

does the table label x and y

OpenStudy (anonymous):

no its and S and P where s is the number of bushels and p is price

OpenStudy (anonymous):

Ok, I'm assuming S for Stock, P for Price. So, it looks like your basic graph of quantity v. price in economics. so P is on the x-axis, S on the y-axis. So, S=mP+b. The only thing is that you then have the change in y as 1/2 and the change in x as 105/100 or 21/20. Giving m=10/21, which makes no sense.

OpenStudy (anonymous):

it states P is on the vertical axis

OpenStudy (anonymous):

meaning the y??

OpenStudy (anonymous):

Ok, so that helps then. Its important to know what is on the x-axis, and what is on the y-axis, but you're right price would be on the y-axis if thats what it says.

OpenStudy (anonymous):

So we really have P=mS+b

OpenStudy (anonymous):

\[\frac{ y2-y1 }{ x2-x1 }=\frac{ 2.4-1.35 }{ 1.5-1 }=\frac{ 1.05 }{ .5 }=2.1\] P=2.1S+b Let's plug in points. Since P = y-axis and S= x-axis, (1, 1.35) is our first point from the table. (1.35)=2.1(1)+b b=1.35-2.1 b=-0.75 Our equation is P=2.1S-0.75 Let's test it: P=2.1S-0.75 with point (2, 3.45) P=2.1S-0.75 P=2.1(2)-0.75 P=4.2-0.75 P=3.45 So the equation works!!!

OpenStudy (anonymous):

phew, thanks a lot buddy! i was thinking to do that but i flipped the S and P around -__- thanks for your patience!

OpenStudy (anonymous):

No problem. Just be sure to keep track of your x and y values and it will make things a lot easier. Even label write out the values as coordinate pairs and label them S and P if you have to. Whatever it takes to keep it in order.

OpenStudy (anonymous):

right. i'll message you again if i need help?

OpenStudy (anonymous):

Sure thing. I try to get on when I can.

OpenStudy (anonymous):

alright